OpenStudy (anonymous):

Select the pair of equations whose graphs are perpendicular. A. 2x – 8y = 9 12x – 3y = 7 B. 3x + 6y = 8 y = 2x – 8 C. y = 2x – 7 x + y = 3 D. 2y = –3x + 5 2x + 3y = 4

OpenStudy (***[isuru]***):

when 2 lines r perpendicular to each other... if the slope of one line is "m" then the slope of the other should be "(-1/m)"

OpenStudy (***[isuru]***):

in the standard form y = mx + c m is the slope... so u have to convert ur answers to standard form and c whether the which pair satisfy the above conditions

OpenStudy (***[isuru]***):

I will show u how to do it for the first answer.. u may follow it to the other ones... Lets consider the 1ST choice 2x - 8y = 9 convert it to the standard form 8y = 2x - 9 y = (2/8)x - 9/8 so the slope of this line is 2/8 the if there is an line perpendicular to this one the slope of that line should be (-8/2) which is(-4) now lets consider the other line 12x - 3y = 7 3y = 12x - 7 y = 4x = 7/4 here the slope is 4 but it should be (-4).. so it;s not this one.... now u have to do the same steps and check the other answers.. can u do that ... i wll help u ??